OpenStudy (anonymous):

Does the integral of pi/(x^2) with limits 1 to infinity converge?

OpenStudy (anonymous):

Ok, so this is what I get, but my calculator actually says that it diverges: \[\int\limits_{1}^{\infty} \Pi/x^2=\lim_{x \rightarrow \infty} -\Pi/x=0-\Pi(-1)=\Pi\]